OpenStudy (anonymous):

Algebra 2 please help

OpenStudy (anonymous):

How much will $850 amount to be in three years if it is invested at 8% interest compounded quarterly for 3 years?

OpenStudy (anonymous):

\[FV = PV \left( 1+\frac{ r }{ n } \right)^{n\cdot t}\] where FV = future value PV = present value r = rate n = number of compunds per yer t = number of years

OpenStudy (anonymous):

what would be the rate? @pgpilot326

OpenStudy (anonymous):

8% but use the decimal representation in the formula

OpenStudy (anonymous):

wow i got 206550 i dont think its right

OpenStudy (anonymous):

\[FV = 850\left( 1+\frac{ .08 }{ 4 } \right)^{4\cdot3}=1078.0055...\]

OpenStudy (anonymous):

i see what i did wrong i put 8 instead of 0.8 thank you

OpenStudy (anonymous):

not 0.8, it's 0.08 = 8/100

OpenStudy (anonymous):

yeah, makes a bit of a difference... 800% vs. 8%